# Brief Theory of Multiqubit Measurement

### 23 Jan 2024 v1

@@ 1 Introduction
Multiqubits are a part a rather broader concept of qudits – quantum objects having Hilbert spaces with finite dimension. Multiqubits have the Hilbert space dimension $N=2^{p}$ corresponding to the dimension of $p$ qubit spaces direct product. Main part of the properties of multiqubits is inherited from qudits. Entanglement of multiqubits can be more complicated than the entanglement of a qubit pair since the qubits can entangle not only in pairs, but in triplets, quadruplets, and even all together [24, 25, 26]. Study of maximally entangled states forms a new topic [27, 28, 29, 30, 7].
@@
In section 4 main aspects of the multiqubit measurement theory are considered. The algorithm for solving the problem on multiqubit state reconstruction by the results of $14^{p}-$ 1 series of measurements for qubit observables comprising a sufficient set of multiqubit incompatible observables is constructed. The properties of the conditional phase portrait and the properties of the measured state resolution of identity transformation to the observables resolution of identity matrices are used to analyse maximally entangled pure and mixed states.
@@ 2 One Particle
    The set of eigenvectors of a Hermitian or a unitary matrix ${X}$:
@@ 2.2.1 Qudit Spaces
Set $\mathcal{B}_{N}$ of the Hilbert space $\mathcal{H}_{N}$ orthonormal bases with typical element $\mathfrak{b}$,
@@
$U\left[A\mapsto\rho\right]_{k,k^{\prime}}=\left\langle k;A|k^{\prime};\rho\right\rangle,\quad U\left[\rho\mapsto A\right]_{k,k^{\prime}}=\left\langle k;\rho|k^{\prime};A\right\rangle,$
@@ 2.2.2 Extension of Pauli Matrices
The coefficients $d_{p}$ characterize the deviation of density matrix from the equilibrium one $\frac{1}{N}\hat{I}$ and determine the expected values of observables represented by matrices $\beta_{p}$
@@ 2.2.3 Qudit Symmetry
The generator $J$ of an arbitrary element $U=\exp\left(iJ\right)$ of this group is a Hermitian matrix with zero trace. Each generator produces an one-parameter Abel subgroup with elements:
@@
The subgroups with one of basis matrices (17) as a generator are periodical ones since the exponent in (22) for such subgroups takes trigonometric form:
@@ 2.2.4 Measures of Reduction
Effect of different angles depends on the measured state since each angle characterizes the reduction of part of the states only. For instance, transformation by the matrix $U=\exp{\left(i\phi\sigma_{x}^{\left\{01\right\}}\right)}$ entangles the measurement of states from the subspace formed by vectors $\left|0\right\rangle$ and $\left|1\right\rangle$, but the states from the subspace with basis $\left\{\left|n\right\rangle,\ n\in\left[2,N-1\right]\right\}$ are measured without changes. Thus, the number of nonzero eigenvalues of transformation matrix generator is an additional characteristic of the reduction of state.
@@ 2.3.1 Qubit spaces
Most elements of the basic space $H=C^{2}$ are redundant, the qubit pure states are represented only with normalized vectors $\left|\psi\right\rangle$ the set of which is equivalent to a two-dimensional sphere $\mathfrak{P}_{2}=S_{2}$. This sphere, it is the Bloch sphere, is the qubit phase space, its points are parametrized by a pair of spherical coordinates $\left\{\vartheta,\varphi\right\}$ or by a three-dimensional unit vector $\vec{n}=\left\{\sin\vartheta\cos\varphi,\sin\vartheta\sin\varphi,\cos\vartheta\right\}$.
@@
Square of generator is proportional to the unit matrix $J^{2}=j^{2}\hat{I}$, so the exponential representation (33) becomes a trigonometric one
@@ 2.3.2 Phase portrait
Figure 1: Examples of qubit phase portraits. Left and centre – phase portraits of qubit pure states with two different positions of state vectors on Bloch sphere. Probability is one in the case of detector direction of which coincides with the direction of the density matrix vector (the lower pole of the phase portrait), and zero in the case of detector with opposite direction (the upper pole of the phase portrait). Right – cross-sections of qubit phase portraits for three listed values of state weight.
@@ 3.1.2 Decomposition of a density matrix
The sets of projectors (43) expand to bases of local algebras. Let us denote
@@
The Hermitian $N\times N$ matrices $B^{\left\{n,n^{\prime}|L\right\}}_{x/y}$, together with projectors $\Pi_{n}^{\left\{L\right\}}$, in amount of $N_{L}^{2}$, form the set $\left\{B_{n}^{\left\{L;N\right\}},n\in\left[1,N_{L}^{2}-1\right]\right\}$ of matrices that is the representation of the basis (17) of Hermitian $N_{L}\times N_{L}$ matrices of qudit with dimension $N_{L}$. Similarly, $N\times N$ matrices $B^{\left\{p,p^{\prime}|S\right\}}_{x/y}$, together with projectors $\Pi_{p}^{\left\{S\right\}}$, form the set $\left\{B_{p}^{\left\{S;N\right\}},p\in\left[1,N_{S}^{2}-1\right]\right\}$ of matrices that is the representation of the basis (17) of Hermitian $N_{S}\times N_{S}$ matrices of qudit with dimension $N_{S}$. The products of this two sets of matrices complete these sets up to a vector basis of a composite qudit algebra of observables. The linear combinations
@@ 3.1.3 Density matrices
The first term is fixed by the condition of normalization of the density matrix, the last term is responsible for the mutual moments of the pairs of local observables, the second and the third term determine the density matrices of the particles
@@ 3.1.4 Measurement of local observables
Local measurement for one of the particles is performed by means of a device represented by a degenerate observable $A^{\left\{dg\right\}}$ of the composite qudit. It is accompanied by transformation of qudit state to subspace corresponding to eigenvalue $A^{\left\{dg\right\}}_{n}$ registered in a specific measurement event number $k$. Transformed density matrix is determined [16] by a normalized result of measured state density matrix $\rho^{\left\{m\right\}}$ wrapping by projectors $\Pi_{n}\left[A^{\left\{dg\right\}}\right]$ to that subspace.
@@ 3.2.1 Stabilizer subgroup
Stabilizer subgroup $U_{\mathfrak{R}}$ of resolution of identity $\mathfrak{R}$ consists of all the elements of the group for which this resolution of identity remains unchanged. Composite qudit resolution of identity (40) leaves unchanged all the elements of the group $SU(N)$ generators of which include in the expression (59) only diagonal elements, i.e. the terms of each sum with indices $n=n^{\prime}\left(N_{L}+1\right)$ and $p=p^{\prime}\left(N_{S}+1\right)$. From the definition of the basis matrices (18) it follows that the projectors of resolutions of identity for the particles are linear combinations of such elements, so the projectors commute with the elements of the stabilizer subgroup; the result of transformation for each projector is the initial projector $U_{\mathfrak{R}}\Pi_{k}U_{\mathfrak{R}}^{+}=\Pi_{k}$. The diagonal elements are included to each of three sums in the expression (59) for the generator of an arbitrary element of the group. The stabilizer is the Abel subgroup with dimension $N-1=N_{L}N_{S}-1$. The conditions
@@ 3.2.3 Entanglement
The entangling transformations have in the third sum (62) terms with products of nondiagonal elements of matrix bases of both particles. The most simple example with generator $\sigma_{x}^{\left\{01;L\right\}}\sigma_{x}^{\left\{01;S\right\}}$ has a following shortcoming: without other generator terms, the local transformation (turn around the both axes $y^{\left\{L\right\}}$ and $y^{\left\{S\right\}}$) turns this product to $\sigma_{z}^{\left\{01;L\right\}}\sigma_{z}^{\left\{01;S\right\}}$, and this one belongs to the stabilizer subgroup. Thus, a linear combination of at least two products is to be considered as a simplest entangling transformation; for instance, a linear combination $a\sigma_{x}^{\left\{01;L\right\}}\sigma_{x}^{\left\{01;S\right\}}+b\sigma_{y}^{\left\{01;L\right\}}\sigma_{y}^{\left\{01;S\right\}}$. This is split into a pair of combinations of ladder matrices (14)
@@ 3.3.3 Phase portrait of a qubit pair
$\begin{array}[]{ll}P^{\left\{0\right\}}\left[\rho\right]\left(\vec{m}^{\left\{0\right\}}\right)=\frac{1}{2}+\frac{1}{2}\vec{m}^{\left\{0\right\}}\cdot\vec{d}^{\left\{0\right\}},&P^{\left\{0\right\}}_{f}\left[\rho\right]\left(\vec{m}^{\left\{0\right\}}|\vec{m}^{\left\{1\right\}}\right)=\frac{1}{2}+\frac{1}{2}\vec{m}^{\left\{0\right\}}\cdot\vec{d}^{\left\{0f\right\}},\\ P^{\left\{1\right\}}\left[\rho\right]\left(\vec{m}^{\left\{1\right\}}\right)=\frac{1}{2}+\frac{1}{2}\vec{m}^{\left\{1\right\}}\cdot\vec{d}^{\left\{1\right\}},&P^{\left\{1\right\}}_{f}\left[\rho\right]\left(\vec{m}^{\left\{1\right\}}|\vec{m}^{\left\{0\right\}}\right)=\frac{1}{2}+\frac{1}{2}\vec{m}^{\left\{1\right\}}\cdot\vec{d}^{\left\{1f\right\}}\\ \end{array}$ (73)
@@
$d_{ab}=kd_{a}^{\left\{0\right\}}d_{b}^{\left\{1\right\}}\ \mapsto\ \vec{d}^{\left\{0f\right\}}=\frac{1+k\vec{m}^{\left\{1\right\}}\cdot\vec{d}^{\left\{1\right\}}}{1+\vec{m}^{\left\{1\right\}}\cdot\vec{d}^{\left\{1\right\}}}\vec{d}^{\left\{0\right\}},\quad\vec{d}^{\left\{1f\right\}}=\frac{1+k\vec{m}^{\left\{0\right\}}\cdot\vec{d}^{\left\{0\right\}}}{1+\vec{m}^{\left\{0\right\}}\cdot\vec{d}^{\left\{0\right\}}}\vec{d}^{\left\{1\right\}}.$ (75)
@@ 3.3.4 Light entanglement
The entangled pure state of a qubit pair can always be reduced to a canonical form that can be obtained from (66)
@@
Entangling transformation of basis for a pair of qubits has a similar form for each vector of entangled basis
@@ 4 Multiqubit
$\mathcal{P}\left(M\right)=\left\{p_{k}=p\left[\rho\right]\left(\left(-1\right)^{b_{0}\left(k\right)}\vec{m}_{0},\ldots,\left(-1\right)^{b_{p-1}\left(k\right)}\vec{m}_{p-1}\right),\ k\in\left[0,2^{p}-1\right]\right\}.$ (81)
@@ 4.1 Multiqubit decomposition
Really, the elements of this subalgebra
$\Sigma_{1}^{\left\{q\right\}}=A^{\left\{q\right\}}+{A^{\left\{q\right\}}}^{+},\quad\Sigma_{2}^{\left\{q\right\}}=iA^{\left\{q\right\}}-i{A^{\left\{q\right\}}}^{+},\quad\Sigma_{3}^{\left\{q\right\}}=\hat{I}-2\Pi^{\left\{q\right\}}$ (84)
@@
$\Sigma_{1}^{\left\{q\right\}}\Sigma_{2}^{\left\{q\right\}}=i\Sigma_{3}^{\left\{q\right\}},\quad\Sigma_{3}^{\left\{q\right\}}\Sigma_{1}^{\left\{q\right\}}=i\Sigma_{2}^{\left\{q\right\}},\quad\Sigma_{2}^{\left\{q\right\}}\Sigma_{3}^{\left\{q\right\}}=i\Sigma_{1}^{\left\{q\right\}}$
@@ 4.2 Density Matrices
Hermitian matrices for multiqubit are represented by combinations of Pauli matrices. It is convenient to group the terms that include one and the same number of Pauli matrices of different particles
@@
$\rho^{\left\{q,r\right\}}=\frac{1}{4}\hat{I}+\frac{1}{4}\sum\limits_{a=1,2,3}d^{\left\{q\right\}}_{a}\Sigma^{\left\{q\right\}}_{a}+\frac{1}{4}\sum\limits_{a=1,2,3}d^{\left\{r\right\}}_{a}\Sigma^{\left\{r\right\}}_{a}+\frac{1}{4}\sum\limits_{a,b=1,2,3}d^{\left\{qr\right\}}_{a,b}\Sigma^{\left\{q\right\}}_{a}\Sigma^{\left\{r\right\}}_{b}.$ (90)

### 24 Feb 2024 v2

@@ 1 Introduction
Multiqubits are a part of a rather broader concept of qudits – quantum objects having Hilbert spaces with finite dimension. Multiqubits have the Hilbert space dimension $N=2^{p}$ corresponding to the dimension of $p$ qubit spaces direct product. Main part of the properties of multiqubits is inherited from qudits. Entanglement of multiqubits can be more complicated than the entanglement of a qubit pair since the qubits can entangle not only in pairs, but in triplets, quadruplets, and even all together [24, 25, 26]. Study of maximally entangled states forms a new topic [27, 28, 29, 30, 7].
@@
In section 4 main aspects of the multiqubit measurement theory are considered. The algorithm for solving the problem on multiqubit state reconstruction by the results of $14^{p}-$ 1 series of measurements for qubit observables comprising a sufficient set of multiqubit incompatible observables is constructed. The properties of the conditional phase portrait and the properties of the measured state resolution of identity transformation to the observable resolution of identity matrices are used to analyse maximally entangled pure and mixed states.
@@ 2 One Particle
    The set of eigenvectors of a Hermitian or a unitary matrix ${X}$,
@@ 2.2.1 Qudit Spaces
The set $\mathcal{B}_{N}$ of the Hilbert space $\mathcal{H}_{N}$ orthonormal bases with typical element $\mathfrak{b}$,
@@
$U\left[A\mapsto\rho\right]_{k,k^{\prime}}=\left\langle k;A|k^{\prime};\rho\right\rangle,\quad U\left[\rho\mapsto A\right]_{k,k^{\prime}}=\left\langle k;\rho|k^{\prime};A\right\rangle$
@@ 2.2.2 Extension of Pauli Matrices
The coefficients $d_{p}$ characterize the deviation of density matrix from the equilibrium one $\frac{1}{N}\hat{I}$ and determine the expected values of observables represented by matrices $\beta_{p}$,
@@ 2.2.3 Qudit Symmetry
The generator $J$ of an arbitrary element $U=\exp\left(iJ\right)$ of this group is a Hermitian matrix with zero trace. Each generator produces an one-parameter Abel subgroup with elements
@@
The subgroups with one of basis matrices (17) as a generator are periodical ones, since the exponent in (22) for such subgroups takes trigonometric form:
@@ 2.2.4 Measures of Reduction
Effect of different angles depends on the measured state, since each angle characterizes the reduction of part of the states only. For instance, transformation by the matrix $U=\exp{\left(i\phi\sigma_{x}^{\left\{01\right\}}\right)}$ entangles the measurement of states from the subspace formed by vectors $\left|0\right\rangle$ and $\left|1\right\rangle$, but the states from the subspace with basis $\left\{\left|n\right\rangle,\ n\in\left[2,N-1\right]\right\}$ are measured without changes. Thus, the number of nonzero eigenvalues of transformation matrix generator is an additional characteristic of the reduction of state.
@@ 2.3.1 Qubit spaces
Most elements of the basic space $H=C^{2}$ are redundant, the qubit pure states are represented only with normalized vectors $\left|\psi\right\rangle$, the set of which is equivalent to a two-dimensional sphere $\mathfrak{P}_{2}=S_{2}$. This sphere, it is the Bloch sphere, is the qubit phase space, its points are parametrized by a pair of spherical coordinates $\left\{\vartheta,\varphi\right\}$ or by a three-dimensional unit vector $\vec{n}=\left\{\sin\vartheta\cos\varphi,\sin\vartheta\sin\varphi,\cos\vartheta\right\}$.
@@
Square of generator is proportional to the unit matrix $J^{2}=j^{2}\hat{I}$, so the exponential representation (33) becomes a trigonometric one,
@@ 2.3.2 Phase portrait
Figure 1: Examples of qubit phase portraits. Left and centre – phase portraits of qubit pure states with two different positions of state vectors on Bloch sphere. Probability is one in the case of detector, direction of which coincides with the direction of the density matrix vector (the lower pole of the phase portrait), and zero in the case of detector with opposite direction (the upper pole of the phase portrait). Right – cross-sections of qubit phase portraits for three listed values of state weight.
@@ 3.1.2 Decomposition of a density matrix
The sets of projectors (43) are expanded to bases of local algebras.
Let us denote
@@
The Hermitian $N\times N$ matrices $B^{\left\{n,n^{\prime}|L\right\}}_{x/y}$, together with projectors $\Pi_{n}^{\left\{L\right\}}$, in amount of $N_{L}^{2}$, form the set $\left\{B_{n}^{\left\{L;N\right\}},n\in\left[1,N_{L}^{2}-1\right]\right\}$ of matrices that is the representation of the basis (17) of Hermitian $N_{L}\times N_{L}$ matrices of qudit with dimension $N_{L}$. Similarly, $N\times N$ matrices $B^{\left\{p,p^{\prime}|S\right\}}_{x/y}$, together with projectors $\Pi_{p}^{\left\{S\right\}}$, form the set $\left\{B_{p}^{\left\{S;N\right\}},p\in\left[1,N_{S}^{2}-1\right]\right\}$ of matrices that is the representation of the basis (17) of Hermitian $N_{S}\times N_{S}$ matrices of qudit with dimension $N_{S}$. The products of these two sets of matrices complete these sets up to a vector basis of a composite qudit algebra of observables. The linear combinations
@@ 3.1.3 Density matrices
The first term is fixed by the condition of normalization of the density matrix, the last term is responsible for the mutual moments of the pairs of local observables, the second and the third terms determine the density matrices of the particles
@@ 3.1.4 Measurement of local observables
Local measurement for one of the particles is performed by means of a device represented by a degenerate observable $A^{\left\{dg\right\}}$ of the composite qudit. It is accompanied by transformation of qudit state to subspace corresponding to eigenvalue $A^{\left\{dg\right\}}_{n}$ registered in a specific measurement event number $k$. Transformed density matrix is determined [16] by a normalized result of measured state density matrix $\rho^{\left\{m\right\}}$ wrapping by projectors $\Pi_{n}\left[A^{\left\{dg\right\}}\right]$ to that subspace,
@@ 3.2.1 Stabilizer subgroup
Stabilizer subgroup $U_{\mathfrak{R}}$ of resolution of identity $\mathfrak{R}$ consists of all the elements of the group for which this resolution of identity remains unchanged. Composite qudit resolution of identity (40) leaves unchanged all the elements of the group $SU(N)$, generators of which include in the expression (59) only diagonal elements, i.e. the terms of each sum with indices $n=n^{\prime}\left(N_{L}+1\right)$ and $p=p^{\prime}\left(N_{S}+1\right)$. From the definition of the basis matrices (18) it follows that the projectors of resolutions of identity for the particles are linear combinations of such elements, so the projectors commute with the elements of the stabilizer subgroup; the result of transformation for each projector is the initial projector $U_{\mathfrak{R}}\Pi_{k}U_{\mathfrak{R}}^{+}=\Pi_{k}$. The diagonal elements are included to each of three sums in the expression (59) for the generator of an arbitrary element of the group. The stabilizer is the Abel subgroup with dimension $N-1=N_{L}N_{S}-1$. The conditions
@@ 3.2.3 Entanglement
The entangling transformations have in the third sum (62) terms with products of nondiagonal elements of matrix bases of both particles. The most simple example with generator $\sigma_{x}^{\left\{01;L\right\}}\sigma_{x}^{\left\{01;S\right\}}$ has a following shortcoming: without other generator terms, the local transformation (turn around the both axes, $y^{\left\{L\right\}}$ and $y^{\left\{S\right\}}$) turns this product to $\sigma_{z}^{\left\{01;L\right\}}\sigma_{z}^{\left\{01;S\right\}}$, and this one belongs to the stabilizer subgroup. Thus, a linear combination of at least two products is to be considered as a simplest entangling transformation; for instance, a linear combination $a\sigma_{x}^{\left\{01;L\right\}}\sigma_{x}^{\left\{01;S\right\}}+b\sigma_{y}^{\left\{01;L\right\}}\sigma_{y}^{\left\{01;S\right\}}$. This is split into a pair of combinations of ladder matrices (14)
@@ 3.3.3 Phase portrait of a qubit pair
$\begin{array}[]{l}P^{\left\{0\right\}}\left[\rho\right]\left(\vec{m}^{\left\{0\right\}}\right)=\frac{1}{2}+\frac{1}{2}\vec{m}^{\left\{0\right\}}\cdot\vec{d}^{\left\{0\right\}},\\ P^{\left\{1\right\}}\left[\rho\right]\left(\vec{m}^{\left\{1\right\}}\right)=\frac{1}{2}+\frac{1}{2}\vec{m}^{\left\{1\right\}}\cdot\vec{d}^{\left\{1\right\}},\\ P^{\left\{0\right\}}_{f}\left[\rho\right]\left(\vec{m}^{\left\{0\right\}}|\vec{m}^{\left\{1\right\}}\right)=\frac{1}{2}+\frac{1}{2}\vec{m}^{\left\{0\right\}}\cdot\vec{d}^{\left\{0f\right\}},\\ P^{\left\{1\right\}}_{f}\left[\rho\right]\left(\vec{m}^{\left\{1\right\}}|\vec{m}^{\left\{0\right\}}\right)=\frac{1}{2}+\frac{1}{2}\vec{m}^{\left\{1\right\}}\cdot\vec{d}^{\left\{1f\right\}}\\ \end{array}$ (73)
@@
$d_{ab}=kd_{a}^{\left\{0\right\}}d_{b}^{\left\{1\right\}}\ \mapsto\quad\begin{array}[]{l}\vec{d}^{\left\{0f\right\}}=\frac{1+k\vec{m}^{\left\{1\right\}}\cdot\vec{d}^{\left\{1\right\}}}{1+\vec{m}^{\left\{1\right\}}\cdot\vec{d}^{\left\{1\right\}}}\vec{d}^{\left\{0\right\}},\\ \vec{d}^{\left\{1f\right\}}=\frac{1+k\vec{m}^{\left\{0\right\}}\cdot\vec{d}^{\left\{0\right\}}}{1+\vec{m}^{\left\{0\right\}}\cdot\vec{d}^{\left\{0\right\}}}\vec{d}^{\left\{1\right\}}.\\ \end{array}$ (75)
@@ 3.3.4 Light entanglement
The entangled pure state of a qubit pair can always be reduced to canonical form that can be obtained from (66)
@@
Entangling transformation of basis for a pair of qubits has similar form for each pair of vectors of entangled basis
@@ 4 Multiqubit
$\begin{array}[]{l}\mathcal{P}\left(M\right)=\\ \left\{p_{k}=p\left[\rho\right]\left(\left(-1\right)^{b_{0}\left(k\right)}\vec{m}_{0},\ldots,\left(-1\right)^{b_{p-1}\left(k\right)}\vec{m}_{p-1}\right),\ k\in\left[0,2^{p}-1\right]\right\}.\\ \end{array}$ (81)
@@ 4.1 Multiqubit decomposition
Really, the elements of this subalgebra,
$\Sigma_{1}^{\left\{q\right\}}=A^{\left\{q\right\}}+{A^{\left\{q\right\}}}^{+},\quad\Sigma_{2}^{\left\{q\right\}}=iA^{\left\{q\right\}}-i{A^{\left\{q\right\}}}^{+},\quad\Sigma_{3}^{\left\{q\right\}}=\hat{I}-2\Pi^{\left\{q\right\}},$ (84)
@@
$\Sigma_{1}^{\left\{q\right\}}\Sigma_{2}^{\left\{q\right\}}=i\Sigma_{3}^{\left\{q\right\}},\quad\Sigma_{3}^{\left\{q\right\}}\Sigma_{1}^{\left\{q\right\}}=i\Sigma_{2}^{\left\{q\right\}},\quad\Sigma_{2}^{\left\{q\right\}}\Sigma_{3}^{\left\{q\right\}}=i\Sigma_{1}^{\left\{q\right\}},$
@@ 4.2 Density Matrices
Hermitian matrices for multiqubit are represented by combinations of Pauli matrices. It is convenient to group the terms that include same numbers of Pauli matrices of different particles:
@@
$\begin{array}[]{l}\rho^{\left\{q,r\right\}}=\\ \frac{1}{4}\hat{I}+\frac{1}{4}\sum\limits_{a=1,2,3}d^{\left\{q\right\}}_{a}\Sigma^{\left\{q\right\}}_{a}+\frac{1}{4}\sum\limits_{a=1,2,3}d^{\left\{r\right\}}_{a}\Sigma^{\left\{r\right\}}_{a}+\frac{1}{4}\sum\limits_{a,b=1,2,3}d^{\left\{qr\right\}}_{a,b}\Sigma^{\left\{q\right\}}_{a}\Sigma^{\left\{r\right\}}_{b}.\end{array}$ (90)